OpenStudy (jango_in_dtown):

If the number x=0.00411522633744855... is written in the standard form as p/q, then show that one of the factors of q-p is 11.

OpenStudy (jango_in_dtown):

@ikram002p

OpenStudy (jango_in_dtown):

hello @Seratul ..

OpenStudy (vanesaretana):

...

OpenStudy (jango_in_dtown):

@vanesaretana can you help me please?

OpenStudy (vanesaretana):

Yes :)

OpenStudy (vanesaretana):

@jango_IN_DTOWN whats the ... options

OpenStudy (jango_in_dtown):

we have to prove that q-p is 11

OpenStudy (jango_in_dtown):

@johnweldon1993

OpenStudy (jango_in_dtown):

@sshayer

OpenStudy (sshayer):

x*1, 00,00,00,00,00,00,000=00,41,15,22,63,37,44,8.55... (1) multiply by 10 x*10,00,00,00,00,00,00,000=41,15,22,63,37,44,85.5... (2) subtract (1) from (2) x(10,00,00,00,00,00,00,000-1,00,00,00,00,00,00,000)=? then find x

OpenStudy (jango_in_dtown):

@sshayer how can we find x. And the method you are saying is used when we know when the digits will repeat, so that the parts after the decimal donot occur after the process . But in this case I dont know how to use it

OpenStudy (sshayer):

from the question it appears 5 is repeating,i may be wrong. But i think so.

OpenStudy (jango_in_dtown):

we cannot make the decimal part zero by the method you are saying.

OpenStudy (sshayer):

if it is repeating then decimal part will become zero.

OpenStudy (phi):

0.00 411 522 633 744 855 if it continues in that pattern, then 966 are the next digits then (maybe??) 10 77 11 88 1299 131010

OpenStudy (jango_in_dtown):

@phi or may be 13110 , if we consider that 11,22,33 is increasing by 11

OpenStudy (jango_in_dtown):

@imqwerty

imqwerty (imqwerty):

hmm is that number even rational

OpenStudy (jango_in_dtown):

I don't have any idea...

imqwerty (imqwerty):

Yeah it is rational Hint to get solution- p/q = 0.00411522633... 1st quation 1000p/q=4.11522633... 2nd equation 2nd equation - 1st equation 1000p/q-p/q =4.11522633.. - 0.00411522.. 999p/q=4.111111111111...

OpenStudy (jango_in_dtown):

great.... :) thank you... :) @imqwerty

imqwerty (imqwerty):

np =)